# Debiased Sample Selection for Combating Noisy Labels

### v1

@@ 4.2 More Analyses
[…]
Ablation study. Our framework mainly contains two modules, the robust network architecture and the mixed data sampling strategy. We conduct ablation studies on CIFAR benchmarks to evaluate the effectiveness of each component (shown in Table 5). First, compared with typical ResNet architecture, our expert-based network indeed makes an obvious contribution to mitigating the noise. The average improvement among five noisy settings roughly reaches 2%. Second, the mixed sampling strategy also promotes the performance of our network. Note that compared with training on randomly sampled batches, training on two weighted sampled batches gains greater generalization. Thus, each module in our learning framework is effective.
Hyperparameter selection. There are two main hyperparameters $m$ and $\beta$ in our framework. We conduct ablation studies to select optimal values for experiments, where $m\in\{2,3,4,5,6,7,8\}$ and $\beta\in\{1.3,1.8,2,3,5,10\}$. The result is shown in Figure 5. For the number of expert layers, we can see that the best test accuracy increases with a larger value of $m$. When $m=4$, the performance reaches the peak and slightly fluctuates as $m$ increases. Meanwhile, the gap between the peak value and the least value is around 0.6% on CIFAR-10 and 0.9% on CIFAR-100, respectively. For the slope parameter in the mapping function, we observe that the performance of our framework is not sensitive to varying values of $\beta$. Therefore, we select $m=4$ and $\beta=3$ for all experiments.
Figure 6: Comparisons of class-level prediction results on CIFAR-100 with 40% symmetric noise. Note that “CE (clean)" denotes training the model on the completely clean set (50k samples).
@@
Class-balanced learning. We visualize the test accuracy in each class to demonstrate that ITEM achieves relatively balanced performance on varying categories while fulfilling greater generalization. We plot comparison results in Figure 6. First, compared with CE (clean), which is trained on 50k clean samples, ITEM achieves almost unbiased prediction results. Second, compared with existing methods, ITEM takes tail classes better into account since two weighted samplers. Therefore, ITEM obtains greater performance in these categories. The result demonstrates our proposal has considerable potential in fulfilling class-balanced learning while tackling label noise.
@@ 5 Conclusion
In this paper, we disclosed the data bias, an implicit bias underlying the sample selection strategy. To solve the training and data bias simultaneously, we proposed ITEM, which integrates a noise-robust multi-experts network and a mixed sampling strategy. First, our structure integrates the classifier with multiple experts and leverages expert layers to conduct selection independently. Compared with the prevailing double-branch network, it exhibits great potential in mitigating the training bias, i.e., the accumulated error. Second, our mixed sampling strategy yields class-aware weights and further conducts weighted sampling. By using the mixup of two weighted mini-batches, we mitigate the issue of sparse representation. The effectiveness of ITEM in real applications is verified on diverse noise types and both synthetic and real-world datasets.
Appendix for Debiased Sample Selection for Combating Noisy Labels
A ITEM with a Semi-Supervised Learning Framework
Considering the large scale of unreliable samples discarded, especially in considerable noise ratios, leveraging these samples through semi-supervised learning is necessary. Therefore, we provide a novel version of ITEM with stronger performance. The algorithm is shown in Algorithm 2. Besides, a strong augmentation CTAugment is used, which refers to [64]. Compared with its original version, the new training framework only adds two operations (shown in Lines 12 and 13). Thus, the code implementation is relatively convenient.
@@ A ITEM with a Semi-Supervised Learning Framework
6:     Split $D_{N}$ into the clean set ${\tilde{D}}_{\rm clean}$ and the unreliable set ${\tilde{D}}_{\rm noise}$ based on $\textsf{Criterion}(\mathbbm{F})$, and discard labels in ${\tilde{D}}_{\rm noise}$.
7:     Calculate weighted vector $\bm{v}=[\bm{w}_{1},...,\bm{w}_{K}]$ on ${\tilde{D}}_{\rm clean}$.
8:     Calculate reversed weighted vector $\bm{\tilde{v}}=[\bm{\tilde{w}}_{1},...,\bm{\tilde{w}}_{K}]$ with Eq. (7).
@@
10:        Randomly sample a mini batch $B_{{\bm{v}}}=\{(\bm{x}_{i},{\tilde{y}}_{i})\}_{i=1}^{b}$ from ${\tilde{D}}_{\rm clean}$ according to $\bm{v}$.
11:        Randomly sample a mini-batch $B_{{\bm{\tilde{v}}}}=\{(\bm{x}^{\prime}_{i},{\tilde{y}^{\prime}}_{i})\}_{i=1}^{b}$ from ${\tilde{D}}_{\rm clean}$ according to $\bm{\tilde{v}}$.
12:        Draw a mini-batch $B=\{(\bm{x}^{\prime\prime}_{i})\}_{i=1}^{2b}$ from ${\tilde{D}}_{\rm noise}$ without labels.
13:        Generate pseudo-labels, then have $B=\{(\bm{x}^{\prime\prime}_{i},{\tilde{y}^{\prime\prime}}_{i})\}_{i=1}^{2b}$, where ${\tilde{y}^{\prime\prime}}_{i}=\mathbbm{E}_{f\sim\mathbbm{F}}[f(\bm{x}^{\prime\prime}_{i})]$.
@@
15:        Update network parameters via minimizing $\mathcal{L}\big{(}f^{c}\big{(}\texttt{MixUp}(B_{{\bm{v}}}\text{ and }B_{{\bm{\tilde{v}}}},B)\big{)}$.
@@
B More experimental results
B.1 Results on CIFAR-100N
We additionally verify the effectiveness of $\ddagger$ ITEM on another human-annotated dataset, CIFAR-100N, whose noise ratio is roughly 40.2%. We adopt the same settings on CIFAR-100 (see Table 8), and the results are shown in Table 6. Our proposal ITEM achieves state-of-the-art performance on both supervised and semi-supervised frameworks.
w/o semi w/ semi
CE 55.50 $\pm$ 0.66% ELR + 66.72 $\pm$ 0.07%
Co-teaching 60.37 $\pm$ 0.27% CORES* 55.72 $\pm$ 0.42%
SOP+ 67.81 $\pm$ 0.23% DivideMix 71.13 $\pm$ 0.48%
ITEM 69.47 $\pm$ 0.18% ITEM* 72.40 $\pm$ 0.19%
Table 6: Test accuracy (%) on the test set of CIFAR-100N.
B.2 Visualization
We verify the effectiveness of $\ddagger$ ITEM on varying noisy conditions and show the following noise-robust results.
Figure 7: Visualization of class-level selection performance of ITEM on CIFAR-100 with 60% symmetric label noise. As the training process proceeds, ITEM gradually improves the F-score in all classes instead of only in head classes.
(a) Training on randomly sampled mini-batch $B$, the best test accuracy is 88.42%.
(b) Training on $B_{\tilde{v}}$ only, the best test accuracy is 90.17%.
(c) Training on $\texttt{MixUp}(B_{\bm{v}},B_{{\bm{\tilde{v}}}})$, the best test accuracy is 93.04%.
Figure 8: Visualizing feature representations on the CIFAR-10 test set by T-SNE. The noise setting is 40% symmetric noisy labels.
(a) CIFAR-10 $w/$ Inst. 40%
(b) CIFAR-100 $w/$ Inst. 40%
Figure 9: Visualization of training curves under three settings. Suppose a small-loss criterion with the loss threshold, we compare our methods (Ours) with the proposed multi-expert network and single network on two noise conditions.
  • robust selection processes. As shown in Figure 7, the total selection performance increases as the training processes, while our framework has not degraded the selection performance on tail classes.
  • robust feature representations. We visualize the learned representation of the test set on CIFAR-10. We compare three training manners and show the results in Figure 8. Compared with training on randomly sampled mini-batch and $B_{\tilde{v}}$, our mixed sampling strategy yields a feature extractor with a strong representation. As shown in Figure 7(c), the distance among varying class clusters is closer, reflecting that the classifier learns more robust decision boundaries.
  • robust training curves. We conduct ablation studies to verify the effectiveness of our proposed robust network architecture. The results on two noisy settings are shown in Figure 9. We can see that compared with a single network (a normal ResNet structure), our expert network achieves more robustness. Meanwhile, the mixed sampling strategy also contributes to improving generalization performance.
Methods Stanford Dogs CUB-200-2011 Avg
Sym 40% Asy 30% Sym 40% Asy 30%
CE 51.42 58.08 64.01 56.02 57.38
Label Smooth 70.22 64.99 54.39 56.80 61.60
Conf. Penalty 68.69 64.50 52.40 54.33 59.98
Co-teaching 49.15 50.50 46.57 50.60 49.21
JoCoR 49.62 53.59 52.64 51.70 51.89
CE + SNSCL 75.27 64.49 68.83 61.48 67.52
$\ddagger$ ITEM 74.92 68.19 70.04 66.58 69.93
Table 8: Detailed settings about training procedure of ITEM.
CIFAR-10 CIFAR-100 CIFAR-10N (all conditions) Clothing-1M
Statistic data
 Class number 10 10 10 14
 Training size 50,000 50,000 50,000 1,000,000
 Testing size 10,000 10,000 10,000 10,000
@@ B.2 Visualization
 Network ResNet-18 ResNet-34 Pre-trained ResNet-50
 Batch size 64 100
 Epoch 200 10
 Warmup epoch 10 30 10 1
@@
 Learning rate (LR) 0.02 0.001
 Weight decay 1e-3 5e-4
 LR scheduler divided by 10 at [100,150]-th epoch divided by 10 at 5-th epoch
@@
B.3 Results on Fine-Grained Noisy Settings

### v3

@@ 4.3 More Analyses
Debias learning for LNL. We visualize the test accuracy in each class to demonstrate that ITEM achieves relatively balanced performance on varying categories while fulfilling greater generalization. We plot comparison results in Figure 5. First, ITEM achieves almost unbiased prediction results compared to CE (clean), which is trained on 50k clean samples. Second, compared with existing methods, ITEM takes tail classes better into account since two weighted samplers. Therefore, ITEM performs better in these categories. Besides, under 20% symmetric label noise, ITEM achieved better performance than the model trained on an absolutely clean set (see the red line vs. the black line). These results demonstrate that our approach has significant potential to address the implicit bias inherent in sample selection frameworks for noisy label learning.
Robustness of MoE structure. Compared with existing robust network, our proposal MoE exhibits better robustness. To clearly show this merit, we evaluate two selection criteria and plot the training curve under three networks in Figure 6. For the small-loss criterion, which easily incurs accumulation errors (see the green line (a)), the performance improvement with the MoE structure is significant. In contrast to the double-branch structure, which experiences slight performance degradation when the learning rate decreases at the 100th epoch (see the orange line in (b)), the learning curve with the MoE structure is more robust and stable. Second, in terms of fluctuation-based selection, the MoE demonstrates greater robustness, more stable training curves, and better generalization performance in noisy settings. In addition, its plug-in plug-out capability further verifies the versatility of the MoE structure.
Table 4: Ablation studies of each component in ITEM with varying noise conditions.
Methods CIFAR-10 CIFAR-100 CIFAR10N CIFAR100N
Sym 40% Sym 40% Worst Fine
w/o MoE Network 90.75 73.01 89.72 68.19
w/o Mixed Sampling 88.77 71.09 87.66 68.24
w/o MixUp 91.19 72.50 89.90 69.10
Ours 92.81 74.90 91.15 69.72
4.4 Ablation Studies
Effect of each component. Our framework mainly contains two modules: the robust network architecture and the mixed data sampling strategy. We conduct ablation studies on CIFAR benchmarks to evaluate the effectiveness of each component from the following two perspectives.
1) Quantitative analysis. We split the main component in ITEM into three parts: robust MoE, Mixed sampling, and the MixUp strategy. By solely removing them from ITEM, we reported the performance of the model in Table 4. Compared with typical ResNet architecture, our expert-based network indeed makes an obvious contribution to mitigating the noise. The average improvement among five settings roughly reaches 2%. Second, the mixed sampling strategy also promotes the performance of our network. Compared with training on randomly sampled batches, training on two weighted sampled batches gains greater generalization.
2) Representation visualization. Considering that training with different sampling leads the model to different representation performance, we compare the learned representation space of the test set of CIFAR-10 and show the results in Figure LABEL:fig:TSNE. We can see that the learned representation of ten classes on the normal sampling $B$ is incompact, while the decision boundary is not well generalized. By contrast, if the proposed reversed tail-focused batch $B_{\tilde{v}}$ is adopted, each cluster is more compact where the distance between each cluster is larger, which verifies the effectiveness of our proposal sampling strategy in mitigating data bias underlying the selection phase. Eventually, by leveraging the Mixup operation, a feature extractor with better representation performance is obtained, as shown in the last plot.
Figure 8: Sensitivity analyses of two hyper-parameters in ITEM. A ResNet-18 is trained on CIFAR-10 with 40% sym. noisy labels.
Sensitivity analyses. ITEM has two main hyper-parameters, $m$ and $\beta$. We conduct ablation studies to select optimal values for experiments. The result is shown in Figure 8. First, the number of expert layers $m$ is the most important parameter in our framework, which decides the overall network structure and computation costs. From the right figure, we can observe that the performance of ITEM gradually increases with the value of $m$ increases. When arriving at a peak, i.e., 93.55% top-1 accuracy at $m=5$, the performance will not increase if the value of $m$ becomes larger. Therefore, we recommend a larger value of $m$ given an unknown dataset. Second, the value of $\beta$ has slightly influenced the performance of our method, where we also prefer a larger value of $\beta$.
5 Related Work
There are three main research lines in sample selection against the incorrect information from noisy labels.
1) Robust selection criteria. Previous sample selection methods normally exploit the memorization effect of DNNs, i.e., DNNs first memorize training data with clean labels and then those with noisy labels [11, 43]. The small-loss criterion [11] is a typical method stemming from the memorization effect, which splits the noisy training set via a loss threshold. The samples with small losses are regarded as clean samples. Some improved criteria [21, 1] based on the small-loss criterion are proposed in later works. In addition to the family of small loss criterion, some methods motivated by prediction fluctuation are designed. These methods [44, 56] observe that the model tends to give inconsistent prediction results for noisy samples and thus filter noisy labels by identifying high-frequency fluctuation samples.
2) Robust network architecture. Easily works based on constructing the noise transition matrix always resort to an additional adaptation layer at the top of the softmax layer [14] or design a new dedicated architecture [10, 49]. Recently, a family of double-branch networks [11, 50, 37, 21] has been proposed, which integrates two networks with the same architecture and selects clean samples for another network. Besides, better performance would be achieved when ensembling prediction results from two networks [11].
3) Robust training procedures. Overall, current methods in sample selection always resort to a self-training manner, i.e., iteratively conducting clean label selection and model retraining. However, the incorrect selection would degrade the subsequent model learning. Recently, some robust training manners were proposed. Concretely, [56] designed a framework based on curriculum learning [5], which starts with learning from clean data and then gradually moves to learn noisy-labeled data with pseudo labels.
Relations to us. The advantages of our proposal compared with previous methods can be divided into the following three parts, which are summarized in Table 5 in Appendix.
  • Flexibility of the network structure. To our knowledge, we are the first to apply a mixture-of-experts (MoE) structure to LNL. Despite the core idea of MoE and double-branch structure being similar (i.e.,ensembling prediction results and the selection result over different views), MoE exhibits more flexibility. Expanding the number of experts is simple and only increases slight parameter counts (0.01 million per expert). By contrast, the double-branch structure is hardly extended to a three- or four-branch structure since the number of learnable parameters.
  • Expansibility of the framework. Compared with previous selection-based methods, our proposal is agnostic to the selection criteria, which can be easily integrated with different selection criteria and improves their performance.
  • Debias learning. Previous mainly focus on the accumulation error, a training bias. In this paper, we empirically show the existence of data bias (i.e., the imbalanced dataset selected by current selection criteria) and design an expert-based sampling strategy, which mitigates this bias by adjusting selection probabilities and increasing the visibility of underrepresented classes.
@@ 6 Conclusion
In this paper, we disclosed the data bias, an implicit bias underlying the sample selection strategy. To solve the training and data bias simultaneously, we proposed ITEM that introduces a noise-robust multi-experts network and a mixed sampling strategy. First, our structure integrates the classifier with multiple experts and leverages expert layers to conduct selection independently. Compared with the prevailing double-branch network, it exhibits great potential in mitigating the training bias, i.e., the accumulated error. Second, our mixed sampling strategy yields class-aware weights and further conducts weighted sampling. The effectiveness of ITEM in real applications is verified on diverse noise types and both synthetic and real-world datasets.
Appendix A Visualization of Data Bias
In this section, we conducted extensive experiments that discovered the existence of the data bias, e.g., the imbalanced subset is selected by current selection-based frameworks. Despite Figure 1 in the section on Introduction, we show more results in the following figure (See Figure 9). We can observe that data bias widely exists in selection-based frameworks no matter the different noise types and different datasets.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 9: Extended results of Figure 1. (a) Small-loss with threshold [11], (b) Small-loss with GMM [21], and (c) Fluctuation-based noise filtering [39].
Table 5: We compare ITEM with various selection-based methods over four perspectives.
Methods Network architecture Selection-criterion specific Training debias Data debias
Co-teaching [11], CoDis [45] Double-branch Small-loss criterion
JoCoR [37], DivideMix [21], ProMix [36] Double-branch Small-loss criterion
SFT [39], Late Stopping [51] Single-branch Fluctuation-based selection
ITEM (ours) Mixture-of-Experts Not specific
Appendix B Algorithm of ITEM
We provided the pseudocode for ITEM in Algorithm 1.
Algorithm 1 ITEM
Input: Noisy training set ${D}_{N}$, training epoch $T$, warmup iterations $T_{w}$, a selection criterion $\textsf{Criterion}(\cdot)$, a classifier network with $m$ experts, and the mapping function $\mathcal{S^{\beta}(\cdot)}$ with a hyper-parameter $\beta$.
2:  for $t$ = 1, …, $T_{w}$ do
3:     Randomly sample $f^{c}$ from $\mathbbm{F}$, and WarmUp $f^{c}$ on $D_{N}$.
4:  end for
5:  for ${t}$ = 1, …, $T$ do
6:     Select the clean set ${D}_{\mathrm{clean}}$ from $D_{N}$ based on $\textsf{Criterion}(\mathbbm{F})$.
7:     Calculate weighted vector $\bm{v}=[{w}_{1},...,{w}_{K}]$ on ${D}_{\mathrm{clean}}$.
8:     Calculate reversed weighted vector $\bm{\tilde{v}}=[{\tilde{w}}_{1},...,{\tilde{w}}_{K}]$ with Equation (7).
9:     while $i$ = 1, …, iterations do
10:        Sample a mini batch $B_{{\bm{v}}}=\{(\bm{x}_{i},{\tilde{y}}_{i})\}_{i=1}^{b}$ from ${D}_{\mathrm{clean}}$ according to $\bm{v}$.
11:        Sample a mini-batch $B_{{\bm{\tilde{v}}}}=\{(\bm{x^{\prime}}_{i},{\tilde{y}^{\prime}}_{i})\}_{i=1}^{b}$ from ${D}_{\mathrm{clean}}$ according to $\bm{\tilde{v}}$.
12:        Sample $f^{c}$ from $\mathbbm{F}$.
13:        Update the network parameters by minimizing the loss in Equation (9).
14:     end while
15:  end for
ITEM with semi-supervised learning. Considering the large scale of unreliable samples discarded, especially in considerable noise ratios, leveraging these samples through semi-supervised learning (SSL) to regularize the model is necessary as well as effective. Thanks to the flexible training framework, any SSL method can be integrated into our proposal ITEM.
Here, we provide a novel version of ITEM with stronger performance. The algorithm is shown in Algorithm 2. Besides, a strong augmentation CTAugment is used, which refers to [32]. Compared with its original version, the new training framework only adds two operations (shown in Lines 12 and 13). Thus, the code implementation is relatively convenient.
@@ Appendix B Algorithm of ITEM
6:     Split $D_{N}$ into the clean set ${D}_{\rm clean}$ and the unreliable set ${\tilde{D}}_{\rm noise}$ based on $\textsf{Criterion}(\mathbbm{F})$, and discard labels in ${\tilde{D}}_{\rm noise}$.
7:     Calculate weighted vector $\bm{v}=[{w}_{1},...,{w}_{K}]$ on ${\tilde{D}}_{clean}$.
8:     Calculate reversed weighted vector $\bm{\tilde{v}}=[{\tilde{w}}_{1},...,{\tilde{w}}_{K}]$ with Eq. (7).
@@
10:        Randomly sample a batch $B_{{\bm{v}}}=\{(\bm{x}_{i},{\tilde{y}}_{i})\}_{i=1}^{b}$ from ${D}_{\rm clean}$ according to $\bm{v}$.
11:        Randomly sample a batch $B_{{\bm{\tilde{v}}}}=\{(\bm{x}^{\prime}_{i},{\tilde{y}^{\prime}}_{i})\}_{i=1}^{b}$ from ${D}_{\rm clean}$ according to $\bm{\tilde{v}}$.
12:        Draw a mini-batch ${\hat{B}}=\{(\bm{x}^{\prime\prime}_{i})\}_{i=1}^{2b}$ from ${D}_{\rm noise}$ without labels.
13:        Generate pseudo-labels for each sample in ${\hat{B}}$, then have ${\hat{B}}=\{(\bm{x}^{\prime\prime}_{i},{\tilde{y}^{\prime\prime}}_{i})\}_{i=1}^{2b}$, where ${\tilde{y}^{\prime\prime}}_{i}=\mathbbm{E}_{f\sim\mathbbm{F}}[f(\bm{x}^{\prime\prime}_{i})]$.
@@
15:        Update network parameters via minimizing $\mathcal{L}\big{(}f^{c}\big{(}\texttt{MixUp}(B_{{\bm{v}}}\text{ and }B_{{\bm{\tilde{v}}}},{\hat{B}})\big{)}$.
@@
Appendix C Experimental Settings
In this section, we reported the statistical data of seven datasets utilized in this paper and the whole experimental setup in Table 6.
C.1 Dataset
CIFAR-10 and CIFAR-100 [17] dataset consists of 60,000 images of 10 and 100 categories. Each image is with a resolution of 32 $\times$ 32.
CIFAR10N and CIFAR100N [38] also consists of 60,000 images. Meanwhile, the noise labels in these two are human-made with varying noise conditions, including worst and random 1&2&3 for CIFAR10N and fine noise for CIFAR100N.
Clothing1M [46] is a large-scale dataset that is collected from real-world online shopping websites. It contains 1 million images of 14 categories whose labels are generated based on tags extracted from the surrounding texts and keywords, causing huge label noise. The estimated percentage of corrupted labels is around 38.46%. A portion of clean data is also included in Clothing1M, which has been divided into the training set (903k images), validation set (14k images), and test set (10k images). We resize all images to $256\times 256$ as in [21] and then random crop to 224 $\times$ 224.
[…]
WebVision [23]. Following [26], we use a mini version of WebVision where only the top 50 classes are utilized. This mini WebVision dataset contains approximately 66 thousand images and the corresponding noise ratio is roughly 20%.
C.2 Other Baselines
On two human-annonated datasets CIFAR10N and CIFAR100N, we compare ITEM with Co-teaching+ [50], Peer Loss [27], CAL [57], DivideMix [21], ELR+ [26], CORES* [8], and DPC [58]. The results are collected from the literature [38]. On three real-world noisy datasets, we compared ITEM with SURE [25], Co-teaching [11], MentorNet [12], JoCoR [37], CDR [43], ELR [26], DivideMix [21], SFT [39], CoDis [45], SURE [25].
Table 6: Experimental settings about training procedure of ITEM.
Datasets CIFAR-10 CIFAR-100 CIFAR-10N CIFAR-100N Clothing-1M Food-101N WebVision
Class number 10 100 10 100 14 101 50
Training size 50,000 50,000 50,000 50,000 1,000,000 310,009 66,000
Testing size 10,000 10,000 10,000 10,000 10,000 25,250 2,500
@@ C.2 Other Baselines
Network ResNet-18 ResNet-34 ResNet-50 InceptionResNetV2
Batch size 64 100 32
Epoch 200 10 100 30
Warmup epoch 10 30 10 30 1 5 3
Learning rate (LR) 0.02 0.001 0.02
Weight decay 1e-3 5e-4
LR scheduler divide 10 at [100,150]th epoch divide 10 at 5th epoch divide 10 at 50th epoch divide 10 at 20th epoch
@@
[…]
Imbalanced ratio $\xi$ 10 100
Noise ratio $\rho$ 10% 20% 30% 40% 50% 10% 20% 30% 40% 50%
CIFAR-10 CE 80.41 75.61 71.94 70.13 63.25 64.41 62.17 52.94 48.11 38.71
Co-teaching [11] 80.30 78.54 68.71 57.10 46.77 55.58 50.29 38.01 30.75 22.85
RoLT [41] 85.68 85.43 83.50 80.92 78.96 73.02 71.20 66.53 57.86 48.98
Sel-CL+ [22] 86.47 85.11 84.41 80.35 77.27 72.31 71.02 65.70 61.37 56.21
CurveNet [13] 87.12 85.02 84.39 82.99 78.57 76.55 74.19 71.90 67.20 64.83
RCAL [54] 88.09 86.46 84.58 83.43 80.80 78.60 75.81 72.76 69.78 65.05
Ours 87.32 86.31 84.91 84.72 81.04 78.39 76.21 74.19 70.99 67.28
CIFAR-100 CE 48.54 43.27 37.43 32.94 26.24 31.81 26.21 21.79 17.91 14.23
Co-teaching [11] 45.61 41.33 36.14 32.08 25.33 30.55 25.67 22.01 16.20 13.45
RoLT [41] 54.11 51.00 47.42 44.63 38.64 35.21 30.97 27.60 24.73 20.14
Sel-CL+ [22] 55.68 53.52 50.92 47.57 44.86 37.45 36.79 35.09 31.96 28.59
CurveNet [13] 55.96 54.60 51.28 48.10 44.61 40.91 39.71 35.90 32.09 29.71
RCAL [54] 57.50 54.85 51.66 48.91 44.36 41.68 39.85 36.57 33.36 30.26
Ours 56.10 55.93 53.07 48.29 45.17 44.00 41.90 40.44 37.28 34.80
Appendix D More Experimental Results
D.1 Results on Imbalanced Noisy Labels
[…]
Dataset construction. Let $T_{ij}(\bm{x})$ denote a probability that the true label $i$ of the instance $\bm{x}$ is corrupted to the noisy label. Therefore, given a noise ratio $rho$, the noise transition matrix satisfies that $T_{ij}(\bm{x})=\rho$ if $i=j$, otherwise $T_{ij}(\bm{x})=\frac{1}{C-1}\rho$, where $C$ denotes the number of classes. We first construct an imbalanced dataset on CIFAR with different imbalanced ratios $\xi=\{10,100\}$ and then corrupt the labels with varying noisy ratio $\rho=\{10\%,...,50\%\}$.
We compare our proposal ITEM with various methods in LNL and imbalanced LNL, including Co-teaching [11], RoLT [41], Sel-CL+ [22], and RCAL [54]. Comparison results are shown in Table 7. Weighted sampling, which mitigates data bias by adjusting selection probabilities and increasing the visibility of underrepresented classes during training, helps ITEM tackle long-tailed classification tasks. On both imbalanced noisy datasets, ITEM exhibits greater generalization under a high noise ratio. For example, ITEM outperforms RCAL by 1.29% and 0.24% under 40% and 50% noise ratio, respectively. On CIFAR-100 with an imbalanced ratio $\xi=100$, ITEM consistently achieves the best generalization performance over other comparison methods. Under an extreme noise rate $\rho=50\%$, ITEM achieves more than 4.5% improvement of top-1 test accuracy. Therefore, we believe that ITEM can also fulfill strong robustness even under imbalanced real-world noisy conditions.
D.2 Visualization of Class-Level Selection
As shown in Figure 10, we further visualize the class-level selection performance of our proposal. We can see that the total selection performance increases as the training progresses. Further, in the tailed classes where the value $F_{k}$ in other selection frameworks would decrease, ITEM gradually improves the F-score in all classes instead of only in head classes, demonstrating that ITEM effectively mitigates the data bias.
Figure 10: Data debias on selection. We plot the class-level selection performance of ITEM on CIFAR-100 with 60% symmetric label noise.
Table 8: Comparison of total training time in hours on CIFAR10N with Worst noise labels with varying experts number.
Metric Single Network Double-branch Network Ours: MoE
CE Co-teaching DivideMix ITEM ITEM ITEM* ITEM*
[11] [21] ($m=3$) ($m=5$) ($m=3$) ($m=5$)
Test accuracy (%) $\uparrow$ 77.69 83.26 92.56 91.15 91.40 93.14 93.32
Times (hours) $\downarrow$ 1.6 4.2 4.4 1.6 1.7 1.8 1.8
Methods Stanford Dogs CUB-200-2011 Avg.
Sym 40% Asy 30% Sym 40% Asy 30%
Cross-Entropy 51.42 58.08 64.01 56.02 57.38
Confidence Penalty [30] 68.69 64.50 52.40 54.33 59.98
Label Smooth [28] 70.22 64.99 54.39 56.80 61.60
Co-teaching [11] 49.15 50.50 46.57 50.60 49.21
JoCoR [37] 49.62 53.59 52.64 51.70 51.89
SNSCL [40] 75.27 64.49 68.83 61.48 67.52
Ours 74.92 68.19 70.04 66.58 69.93
D.3 Training Time Analysis
In Table 8, we compare the training times of ITEM and (semi) ITEM* with three typical methods, using a single Nvidia 4090 GPU. We can observe that the training time of these methods based on double-brach networks (such as Co-teaching and DivideMix) is always twice as slow as training on a single network. However, in our proposed MoE structure, increasing the number of experts number will not significantly increase the computation cost. Besides, integrating ITEM with a semi-supervised learning framework can bring considerable performance improvement with fewer extra costs.
D.4 Results on Fine-Grained Noisy Labels